OpenStudy (anonymous):

X^3/2=27, find value of x

sam (.sam.):

My guess is \[x^{3/2}=27\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=27^2/3 x=((27)^1/3)^2=(3)^2=9

OpenStudy (anonymous):

27=3^3 27^1/3=(3^3)^1/3=3

sam (.sam.):

Square both sides, you'll get \[x^3=27^2\] \[x^3=729\] Then the cubed of 9 is 729, so rewrite it as \[x^3=9^3\] Then 3 cancels out from the exponents

OpenStudy (anonymous):

answer is correct thanks

OpenStudy (anonymous):

means 9 is answer